write the slope-intersecy form of an equation of the line that passes through the given point and is parallel to the graph of each equation

(3,2),y=x+5

Answer by stanbon(75887)

write the slope-intersecy form of an equation of the line that passes through the given point and is parallel to the graph of each equation
(3,2),y=x+5
-----
The given line has slope = 1
---
Every line parallel to it must have slope = 1
---
Find equation of line with m=1 and y=2 when x = 3.
-----
Form:
y = mx + b
Substitute and solve for "b":
2 = 1*3+b
b = -1
----
Equation:
y = x-1
============
Cheers,
Stan H.